OpenStudy (anonymous):

Follow the process of completing the square to solve x2 - 10x + 8 = 0. What is the value of the constant that will be isolated on the right side of the equation in step 3?

OpenStudy (anonymous):

if you want solve this equation x1 = 5+(17)^0.5 x2 = 5-(17)^0.5

OpenStudy (anonymous):

add and subtract (co=efficient of x/2)^2 inthis case (-10/2)^2 i.e.,25